OpenStudy (anonymous):

Differentiate the function g(x) = ln [x*sqrt(x^2 - 1)]

OpenStudy (anonymous):

\[\ln (x \sqrt{x ^{2}-1}\]

OpenStudy (anonymous):

\[g(x)=\ln[x*\sqrt(x^2-1)]=lnx+\ln \sqrt{x ^{2}-1}=lnx+\frac{ 1 }{2 }\ln \left( x ^{2} -1\right)\] g'(x)=calculate

OpenStudy (anonymous):

thanks!